Circular No. 8409

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004dy, 2004dz, 2004ea, 2004ee, 2004ej, 2004eo G. Folatelli, M. Hamuy and N. Morrell, Carnegie Supernova Project,

report that CCD spectra (range 385–910 nm), obtained on Sept. 14 UT with the Las Campanas du Pont 2.5-m telescope (+ WFCCD spectrograph), reveal that SN 2004dz (IAUC 8395), SN 2004ee (IAUC 8399) and SN 2004ea (IAUC 8396) are evolved type-Ia supernovae, a few months past maximum. They confirm the report on IAUC 8404 that SN 2004dy is a type-II supernova with broad Hα, but they detect a strong He I 5876-nm emission atop a blue continuum, and this suggests a young supernova. SN 2004ej (IAUC 8405) is of type II with a prominent Hα P-Cyg profile, the velocity inferred from the absorption minimum of Hβ being 6400 km/s (assuming the 2723 km/s recession velocity from NED database); both the expansion velocity and the presence of Fe II 516.9-nm suggest that this supernova exploded 20–30 days ago.

Morrell, Folatelli, and Hamuy add that two spectra of SN 2004eo (cf. IAUC 8406), taken on Sept. 19.02 (range 455–960 nm) and Sept. 20.01 UT (range 385–920 nm), reveal that this is a type-Ia supernova displaying the characteristic Si II (635.5, 413.0, and 597.2 nm) and S II (545.4 and 564.0 nm) features. The minimum of the Si II 635.5-nm line yields expansion velocities of 12100 and 10000 km/s, respectively, on the two nights, assuming the NED host-galaxy recession velocity of 4707 km/s. Photometric observations obtained with the 1-m Swope telescope by S. Gonzalez and W. Krzeminski yield the following B-band magnitudes: Sept. 18.00, 17.59; 18.99, 17.13; 19.99, 16.74; 21.04, 16.47; 22.05, 16.21. These suggest that SN 2004eo will reach maximum light in the next 3 days or so.

SUPERNOVAE 2004ed AND 2004ek Further to IAUC 8408, M. Modjaz et al. report that a spectrum of

SN 2004ed (cf. IAUC 8397), obtained by Calkins on Sept. 10.17, shows it to be a late-time type-II supernova with strong Hα emission, an Hβ P-Cyg profile and other elements in absorption. Adopting the NED recession velocity of 7500 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 6200 km/s. A spectrum of SN 2004ek (cf. IAUC 8405), obtained on Sept. 14.47, shows it to be most probably a young type-II supernova. The spectrum consists of a featureless and blue continuum and is similar to an early spectrum of SN 1993J (Matheson et al. 2000, A.J. 120, 1487).

2004 September 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green